OpenStudy (anonymous):

Please help Write the slope-intercept form of the line that passes through the point (1, 0) and is parallel to x - y = 7

OpenStudy (anonymous):

slope intercept form is y=mx+b now you need to isolate the y value to one side

OpenStudy (anonymous):

If the line is parallel to that line, than the slopes must be the same. Then use the point and solve for b (in y = mx + b)

OpenStudy (anonymous):

so how do i do it? i am kinda confused

OpenStudy (anonymous):

y = x - 7 : Slope = 1 0 = (1)(1) + b b = -1 y = x - 1

OpenStudy (anonymous):

Thank you everyone.